# Balanced Mixed-Type Tabular Data Synthesis with Diffusion Models

### v1

@@ 4.2 Backbone
In this research, we model mixed-type tabular data in latent space following a similar setup in (Rombach et al., 2022). The tabular data are encoded into latent space and decoded back with multilayer perceptrons. We choose U-Net (Ronneberger et al., 2015) with transformers as a posterior estimator to predict $\bar{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c},\mathbf{S})$ and $\bar{\mathbf{x}}_{0}(\mathbf{z}_{t},\mathbf{c},\mathbf{S})$. This choice is based on the hypothesis that the U-Net, augmented by the contextual understanding capabilities of transformers, can effectively manage the heterogeneous nature of tabular data. This integration aims to leverage the U-Net’s ability to capture spatial correlation and the transformers’ strengths in sequence modeling, thereby offering a novel approach to latent space modeling of tabular data.
@@ 4.3 Balanced Sampling
After training diffusion models on tabular data paired with labels $\mathbf{c}$ and a set of sensitive features $\mathbf{S}$, we apply multivariate latent guidance for conditional tabular data synthesis. We customize the marginal distributions of the provided labels and sensitive features in the sampling phase. Fair synthetic tabular data are typically anticipated to exhibit the same label distribution as the real data. These data are expected to mirror the label distribution of the real data and exhibit a balanced distribution of sensitive features, which addresses the common issue of label imbalance in real-world datasets. To achieve this, we can compute the empirical distribution of the labels in the actual data and use it together with uniform categorical distributions for each sensitive feature to guide the sampling process. By doing this, we generate balanced synthetic data regarding the sensitive features, and the synthetic data have the same label distribution as the actual data.
@@ 5 Experiments
We first introduce experiment datasets, data processing, baselines, and how we assess synthetic tabular data. Then we present computational results on experimental datasets.
@@ 5.1 Datasets
We evaluate the effectiveness and fairness of our proposed method using seven tabular datasets for classification tasks. These datasets contain numerical and categorical features, including some sensitive attributes with observed class imbalances. We provide a summary of the datasets, including their corresponding sensitive attributes in Table 1, and offer further details, such as data sources, sensitive class label distributions, etc., in Appendix A. Common sensitive attributes such as sex and race can be found in data sets like UCI Adult, Cardio, Credit Card, Depression, KDD Census, and Law School. Additionally, other sensitive attributes such as marital status and education are present in the Bank Marketing and Credit Card datasets. Each data set is divided into training, validation, and test sets using a 50/25/25 split. Following the training of our diffusion models with multivariate guidance, we proceed to generate synthetic data that adheres to the original label distribution while incorporating a uniform distribution for sensitive features.
Table 1: Experimental datasets
Abbreviation Dataset Size Sensitive Target
AD UCI Adult 48842 Sex, Race Income
BA Bank Marketing 45221 Marital Subscription
CA Cardio 70000 Sex Diagnosis
CR Credit Card 30000 Sex, Education Payment
DE Depression 11000 Sex, Race Diagnosis
KD KDD Census 199523 Sex, Race Income
LA Law School 20800 Sex, Race Admission
@@ 5.2 Data Processing and Baselines
Data Processing. Following the data processing steps in (Kotelnikov et al., 2023), we split the numerical and categorical features for each dataset and pre-process them separately. Numerical features are converted through quantile transformation, a non-parametric technique transforming variables into a standard Gaussian distribution based on quantiles, thus normalizing the data and enhancing the robustness to outliers. In preparation for the Markov transition process in the latent diffusion model, we encode categorical features into one-hot vectors and concatenate the resulting vectors with normalized numerical features.
Baselines. In assessing the efficacy of our approach for various datasets, we conduct a comparative analysis against several baselines for tabular data synthesis. Within learning-based methods, we categorize them into diffusion models, generative adversarial networks, and variational autoencoders, selecting the top-performing model from each category with publicly available code. In addition to learning-based techniques, we incorporate SMOTE, a robust and traditional method for tabular data synthesis, into our set of baselines. Specifically, the selected baseline methods for comparison are described in detail as follows.
  • TabDDPM (Kotelnikov et al., 2023): a recent diffusion-based model that has demonstrated to surpass existing methods for synthesizing tabular data. Leveraging multilayer perceptrons as its backbone, TabDDPM excels in learning unbalanced distributions present in the training data.
  • CTGAN (Xu et al., 2019): a tabular data synthesis approach based on generative adversarial networks, providing publicly available code that is accessible and user-friendly in (Patki et al., 2016).
  • TVAE (Xu et al., 2019): a state-of-the-art variational autoencoder specifically designed to generate tabular data with a user-friendly implementation in (Patki et al., 2016).
@@
  • SMOTE (Chawla et al., 2002): a robust interpolation-based method that creates synthetic data points by combining a real data point with its $k$ nearest neighbors from the dataset. In our application, we leverage SMOTE to generate additional tabular data samples, employing interpolation among existing samples with the same label. The implementation of SMOTE is available in (Lemaître et al., 2017).
@@ 5.3 Assessing Synthetic Tabular Data
Model-Based Evaluation. The computation procedure of model-based evaluation of synthetic tabular data in classification tasks involves the following steps.
  1. 1.
    Train a generative model on the original training set.
  2. 2.
    Generate synthetic data with the generative model.
  3. 3.
    Train classifiers on the synthetic data.
  4. 4.
    Evaluate classifiers on the original validation data.
@@
Specifically, when assessing model accuracy on the original validation data, it is referred to as machine learning efficiency as used in (Choi et al., 2017). Beyond accuracy assessments, the evaluation of synthetic data can extend to fairness scores of classifiers, providing insights into the fairness of the generated data. We evaluate the fairness of machine learning models trained on the synthetic data, focusing on two key aspects: equal allocation and equal performance (Agarwal et al., 2018).
@@
  • Equal allocation, a fundamental principle in model-based fairness evaluations, entails the idea that a model should distribute resources or opportunities proportionally among different groups, irrespective of their affiliation with any privileged group. Equal allocation is quantified using the demographic parity ratio. This ratio reveals the extent of the imbalance by comparing the lowest and highest selection rates (the proportion of examples predicted as positive) between groups.
@@
  • Equal performance is grounded in the principle that a fair model should exhibit consistent performance across all groups. This entails maintaining the same precision level for each group. The equalized odds ratio, which represents the smaller of two ratios comparing true and false positive rates between groups, serves as a metric for assessing performance fairness in this context.
@@
The evaluation of the demographic parity ratio and equalized odds ratio depends on the selection of sensitive features. When evaluating fairness scores in our experiments, we choose the first sensitive feature specified in Table 1. Both the demographic parity ratio and equalized odds ratio are within the range $[0,1]$. The demographic parity ratio highlights the importance of distributing resources equally among different groups, whereas the equalized odds ratio prioritizes impartial decision-making within each specific group. Higher values in either metric indicate progress toward fairness. However, attaining perfect fairness, represented by a score of $1$, may not always be feasible or appropriate depending on the specific context. Our proposed method is expected to have enhanced fairness scores compared to baseline models, and we interpret the results using SHAP feature importance (Lundberg & Lee, 2017).
Data-Based Evaluation. The data-based evaluation directly compares synthetic and real tabular data. To address the privacy concerns when sharing tabular data publicly, the synthetic tabular data is expected not to be a “copy” of the original data. The originality of synthetic data is quantified by utilizing a nearest-neighbor distance metric suggested by (Park et al., 2018), known as the “distance to the closest record” (DCR), calculated as the Euclidean distance between each synthetic sample and its nearest real data point. Larger DCR values indicate that the synthetic data has a greater dissimilarity to the original data. In addition, when evaluating the fairness of synthetic tabular data with a data-based method, we examine the imbalance in class characteristics within sensitive features such as sex or race.

### v3

@@ 4.2 Backbone
In this research, we model mixed-type tabular data in latent space following a similar setup in (Rombach et al., 2022). The tabular data are encoded into latent space and decoded back with multilayer perceptrons. We choose U-Net (Ronneberger et al., 2015) with transformers as a posterior estimator to predict $\bar{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c},\mathbf{S})$ and $\bar{\mathbf{x}}_{0}(\mathbf{z}_{t},\mathbf{c},\mathbf{S})$. This choice is based on the hypothesis that the U-Net, augmented by the contextual understanding capabilities of transformers, can effectively manage the heterogeneous nature of tabular data. This integration leverages the U-Net’s ability to capture spatial correlation and the transformers’ strengths in sequence modeling, thereby offering a novel approach to latent space modeling of tabular data.
@@ 4.3 Balanced Sampling
After training diffusion models to approximate the distribution of the training data conditioned on labels $\mathbf{c}$ and sensitive features $\mathbf{S}$, we can generate samples from this conditional distribution and align them with any desired target distribution by adjusting the conditioning variables. Specifically, during sampling, the denoising neural network utilizes guidance from the target label and sensitive attributes. The joint distribution of the target label and sensitive attributes can be customized. In our bias mitigation framework, we expect the synthetic data to preserve the same label distribution as the real data, while making the sensitive attributes nearly independent of the target label. To achieve this, we first compute the empirical label distribution from the real data and apply uniform categorical distributions for each sensitive feature. This results in a balanced joint distribution of the target label and sensitive features, which serves as input for the conditional generation in our diffusion model. Then our diffusion model iteratively denoises the Gaussian noise with label and sensitive guidance to generate fair tabular data.
@@ 5 Experiments
In this section, we first introduce experimental datasets, data processing, baselines, and how we assess synthetic tabular data. Then we present computational results on experimental datasets. The training settings are presented in Appendix B.
@@ 5.1 Datasets
We evaluate the effectiveness and fairness of our proposed method using three binary classification tabular datasets: Adult (Kohavi et al., 1996), Bank Marketing (Moro et al., 2014), and COMPAS (Larson et al., 2016). These datasets contain numerical and categorical features, including some sensitive attributes with observed class imbalances or strong correlations to the target variable. We provide a summary of the datasets, including their corresponding sensitive attributes in Table 1, and offer further details in Appendix A. Common sensitive attributes such as sex and race can be found in Adult and COMPAS. As mentioned in (Rajabi & Garibay, 2022), the age group in the Bank Marketing dataset is considered sensitive, with individuals below 25 classified as “young” and those above 25 classified as “old”. In our experiments, each data set is divided into training, validation, and test sets using a 50/25/25 split. Following the training of our diffusion models with multivariate guidance, we proceed to generate synthetic data that adheres to the original label distribution while incorporating a uniform distribution for sensitive features.
Table 1: Experimental datasets.
Dataset Train Validation Test Sensitive Target
Adult 22611 11305 11306 Sex & Race Income
Bank 22605 11303 11303 Age Subscription
COMPAS 8322 4161 4161 Sex & Race Recidivism
@@ 5.2.1 Data Processing
Following the data processing steps in Kotelnikov et al. (2023), we split the numerical and categorical features for each dataset and pre-process them separately. Numerical features are converted through quantile transformation, a non-parametric technique transforming variables into a standard Gaussian distribution based on quantiles, thus normalizing the data and enhancing the robustness to outliers. In preparation for the Markov transition process in the latent diffusion model, we encode categorical features into one-hot vectors and concatenate the resulting vectors with normalized numerical features.
5.2.2 Baselines
To evaluate the effectiveness of our approach across different datasets, we perform a comparative analysis against state-of-the-art (SOTA) baseline models that have publicly available code for tabular data synthesis. Additionally, we compare our method with fair tabular generative models based on GAN and SMOTE. The specific baseline methods chosen for comparison are detailed below.
  • Fair Class Balancing (FairCB) (Yan et al., 2020): a SMOTE-based oversampling method to adjust class distributions in the training data to address class imbalances.
  • TabFairGAN (FairTGAN) (Rajabi & Garibay, 2022): a fairness-aware GAN tailored for tabular data incorporates a fair discriminator to ensure equality in the joint probability distribution of sensitive features and labels.
  • CoDi (Lee et al., 2023): a recent diffusion-based approach that combines two diffusion models: one for continuous variables and another for discrete variables. These models are trained separately to effectively learn the distributions of their respective data types. However, they are also linked together by conditioning each model’s sample generation on the other model’s samples at each time step, which enhances the learning of correlations between continuous and discrete variables. Additionally, to further strengthen the connection between the two models, contrastive learning is applied to both diffusion models. Negative samples are created by breaking the inter-correlation between the continuous and discrete variable sets, thus reinforcing the models’ ability to learn these relationships.
  • GReaT (Borisov et al., 2022): utilizes pre-trained transformer-decoder language models (LLMs) to generate synthetic tabular data through a two-stage process. In the first stage, it transforms each row of the real tabular dataset into a textual representation by representing each feature with its name and value in a subject-predicate-object structure. The transformation of these features is then concatenated in a random order to form the feature vector of the corresponding row. The random ordering enables the LLM to be fine-tuned on samples without depending on the sequence of features, allowing for arbitrary conditioning during tabular data generation. This textual representation of the tabular dataset is used to fine-tune the pre-trained LLM. In the second stage, the fine-tuned LLM generates synthetic data by initializing it with any combination of feature names or feature name-value pairs, allowing the model to complete the remaining feature vector in its textual representation, which can then be converted back into tabular format.
@@ 5.2.2 Baselines
  • SMOTE (Chawla et al., 2002): a robust interpolation-based method that creates synthetic data points by combining a real data point with its $k$ nearest neighbors from the dataset. In our application, we leverage SMOTE to generate additional tabular data samples, employing interpolation among existing samples with the same label. The implementation of SMOTE is available in Lemaître et al. (2017).
  • STaSy (Kim et al., 2022): a Score-based Tabular Data Synthesis (STaSy) method that adopts the score-based generative modeling approach. It introduces a self-paced learning method that begins by training on simpler records with a denoising score matching loss below a certain threshold. As training progresses and the model becomes more robust, progressively more challenging records are included in the training. Additionally, STaSy includes a fine-tuning approach that adjusts the model parameters by iteratively retraining it using the log probabilities of the generated samples.
  • TabDDPM (Kotelnikov et al., 2023): a recent diffusion-based model that has demonstrated to surpass existing methods for synthesizing tabular data. Leveraging multilayer perceptrons as its backbone, TabDDPM excels in learning unbalanced distributions in the training data.
  • TabSyn (Zhang et al., 2023): a SOTA method that utilizes a latent diffusion model to generate tabular data containing both numerical and categorical features. First, a VAE model, specifically designed for tabular-structured data, converts raw tabular data with both numerical and categorical features into embeddings in a regularized latent space that has a distribution close to the standard normal distribution. Following this, a score-based diffusion model is trained in the embedding space to capture the distribution of these latent embeddings. The diffusion model adds Gaussian noise that increases linearly over time in the forward process, which minimizes approximation errors in the reverse process, thus allowing for synthetic data to be generated in significantly fewer reverse steps.
@@ 5.3 Assessing Synthetic Tabular Data
5.3.1 Data-Based Evaluation
We evaluate synthetic tabular data by directly comparing it with real data, with a focus on column-wise density estimation and pair-wise column correlations. To address privacy concerns when sharing tabular data publicly, synthetic data should not replicate the original data. As suggested by (Zhang et al., 2023), we compute “distance to the closest record” (DCR) scores to ensure this. Furthermore, when assessing the fairness of synthetic tabular data using data-based methods, we analyze the class imbalances within sensitive features like sex or race, as well as the joint distribution with the target label.
5.3.2 Model-Based Evaluation
The model-based evaluation process for synthetic tabular data in classification tasks involves the following steps: first, training a generative model on the original training set; next, generating synthetic data using the trained generative model; then, training classifiers on the synthetic data; and finally, evaluating these classifiers on the original test data. Specifically, when assessing model accuracy on the original test data, it is referred to as machine learning efficiency as used in (Choi et al., 2017).
@@ 5.3.2 Model-Based Evaluation
Beyond accuracy assessments, the evaluation of synthetic data can extend to fairness scores of classifiers, providing insights into the fairness of the generated data. We evaluate the fairness of machine learning models trained on the synthetic data, focusing on two key aspects: equal allocation and equal performance (Agarwal et al., 2018).
@@
  • Equal allocation, a fundamental principle in model-based fairness evaluations, entails the idea that a model should distribute resources or opportunities proportionally among different groups, irrespective of their affiliation with any privileged group. Equal allocation is quantified using the demographic parity ratio (DPR). This ratio reveals the extent of the imbalance by comparing the lowest and highest selection rates (the proportion of examples predicted as positive) between groups.
@@
  • Equal performance is grounded in the principle that a fair model should exhibit consistent performance across all groups. This entails maintaining the same precision level for each group. The equalized odds ratio (EOR), which represents the smaller of two ratios comparing true and false positive rates between groups, serves as a metric for assessing performance fairness in this context.
@@
Both the demographic parity ratio and equalized odds ratio are within the range $[0,1]$. The demographic parity ratio highlights the importance of distributing resources equally among different groups, whereas the equalized odds ratio prioritizes impartial decision-making within each specific group. Higher values in either metric indicate progress toward fairness. In our experiments, when evaluating the fairness scores of classifiers as shown in Table 4, we evaluate the fairness scores of classifiers by selecting sex as the sensitive attribute for the Adult and COMPAS datasets, and age group for the Bank dataset.